### v2

# Exploring singularities in data with the graph Laplacian: An explicit approach

Abstract.
We develop theory and methods that use the graph Laplacian to analyze the geometry of the underlying manifold of datasets. Our theory provides theoretical guarantees and explicit bounds on the functional forms of the graph Laplacian when it acts on functions defined close to singularities of the underlying manifold. We use these explicit bounds to develop tests for singularities and propose methods that can be used to estimate geometric properties of singularities in the datasets.
Key words and phrases:
Graph Laplacian, geometry, singularities
1991 Mathematics Subject Classification:
Primary 58K99; Secondary 68R99, 60B99.
1. Introduction
High dimensional data is common in many research problems across academic fields, and such data can often be represented as points collected or sampled from $\mathbb{R}^{N}$. A common assumption is that the dataset $X=\{X_{i}\}_{i}^{n}$ $\subset\mathbb{R}^{N}$ lies on a lower-dimensional set $\Omega$ and is in fact a sample from some probability distribution over $\Omega$. A further assumption, that makes our model of data more tractable, is that $\Omega$ can be represented as the union of several well-behaved manifolds, i.e. $\Omega=\cup_{i}\Omega_{i}$. Here, each $\Omega_{i}$ could represent a different class in a classification problem: For instance, if a dataset contains two classes, $i$ and $j$, class $i$ might be contained in $\Omega_{i}$ and class $j$ in $\Omega_{j}$, with the two classes potentially being disjoint. However, classification is not always so clear-cut: For instance, in the MNIST dataset, handwritten digits of "$1"\in\Omega_{1}$ and "$7"\in\Omega_{7}$ can appear very similar, suggesting that $\Omega_{1}\cap\Omega_{7}\neq\emptyset$. Therefore, understanding geometric situations such as intersections is of interest in classification problems.
In our manifold model of data, an intersection between two different manifolds $\Omega_{i},\Omega_{j}$ is either represented just as such, or it can be viewed as a singularity if we consider $\Omega=\Omega_{i}\cup\Omega_{j}$ as a single manifold. Other regions in $\Omega$ that can be viewed as singular, such as boundaries and edges, may also be of interest as they can signify important features in the data. Regions that are part of just one manifold $\Omega_{i}$, and in its interior, we consider as non-singular.
To study such singularities, we use the graph Laplacian $L_{n,t}$. This operator, which depends on the number of data points $n$ and a parameter $t$, can act on functions defined on the dataset $X$. In non-singular regions, as $n$ tends to infinity and $t$ tends to 0, $L_{n,t}$ converges to the Laplace-Beltrami operator [1]. In this work, we primarily study the behavior of $x\to L_{n,t}f(x)$ for functions $f$, when $x$ is close to singular points.
1.1. Motivation
As our data lives in the ambient space and we a-priori do not have information about the underlying manifold, we will explore how the graph Laplacian applied to the function $f(x)=v\cdot x$ for a unit vector $v$, defined on the data-set in the ambient space, which is just a vector $\mathbf{f}=(v\cdot X_{1},\ldots,v\cdot X_{n})\in\mathbb{R}^{n}$. Our contribution of this paper is that we calculate explicitly how the graph Laplacian acts on such linear functions $f$, which can then be used for estimation purposes.
The specific choice of $f$ is motivated by the fact that since $L_{n,t}$ converges to the Laplace-Beltrami operator, in the interior of $\Omega$, we expect that for $f$ as above $L_{n,t}f(x)\approx 0$. However, for singular points like intersections, the limit operator is of first order [2], and $L_{n,t}f(x)\neq 0$, which can be seen in Fig. 1.
Our results show how $x\to L_{t}f(x)$ and, through a finite-sample bound, how $x\to L_{n,t}f(x)$ behaves. More specifically, given $x_{0}\in\Omega_{i}$ near some singularity, and $x$ in the ball $B_{R}(x_{0})$, including the case when $x\not\in\Omega_{i}$, we show how the function $x\to L_{n,t}f(x)$ deviates from being constantly 0 and has specific functional forms. These forms depend on the type of singularity.
1.2. Overview of results
First, in Section 4.1 we consider the case that $\Omega$ is flat manifold of dimension $d$, and where we have a geometric situation similar to Fig. 3. This is mainly motivated by the example in Section 2.1.
To set up the results, we start with an $x_{0}\in\Omega$ which is assumed to be a flat manifold, and let $x\in B_{R}(x_{0})$, where $R=\sqrt{t}r_{0}>0$, and use $\hat{x}$ to denote the projection of $x$ to the tangent space of $\Omega$. We also define $v_{n,\Omega}$ as the projection of $v$ onto $x-\hat{x}$, and $v_{n,\partial\Omega}$ is the projection of $v$ onto the outwards normal of $\partial\Omega$, which is assumed to be flat close to $x$. Then we show the following:
  • In Theorem 2, we let $\|x-x_{0}\|=r\sqrt{t}$ and $\theta$ is the angle between vectors $x-x_{0}$ and $\hat{x}-x_{0}$. If $x$ is not close to $\partial\Omega$, then
    $L_{t}f(x)=A(x)t^{\frac{d+1}{2}}v_{n,\Omega}\sin(\theta)re^{-\sin^{2}(\theta)r^{2}}+t^{\frac{d+1}{2}}B(x)e^{-r_{0}^{2}}.$
    The function $A$ is close to being constantly equal to $\pi^{d/2}$, and $B$ is uniformly bounded. Note that for small $t$ we have $r_{0}$ can be taken large and thus the last term is very small. Both functions have explicit bounds.
  • Theorem 3 shows what happens when $x$ is close to $\partial\Omega$:
    $L_{t}f(x)$ $=\widehat{A}_{1}(x)t^{\frac{d+1}{2}}v_{n,\Omega}\sin(\theta)re^{-\sin^{2}(\theta)r^{2}}$
    $\qquad+\widehat{A}_{2}(x)t^{\frac{d}{2}}v_{n,\partial\Omega}e^{-\sin^{2}(\theta)r^{2}}\quad+t^{\frac{d+1}{2}}B(x)e^{-r_{0}^{2}},$
    where functions $\widehat{A}_{1},\widehat{A}_{2}$ and $B$ have explicitly computable bounds.
In Section 4.2 and Section 4.3 we prove more general results:
  • In Theorem 4 we relax the conditions on $\Omega$, considering non-flat manifolds, and prove a weaker version of Theorem 2.
  • In Theorem 5 we relax the conditions further, and allow for noise when sampling from $\Omega$.
In Section 5 we show how the results can be used to construct hypothesis tests for intersections in data. We test this method in Section 6.1. We also show how such hypothesis tests can be used to detect singularities in zero sets of neural networks in Section 6.1.1.
Finally, in Section 6.2 we propose methods to find intersections in data and estimate the angle of such intersections, which are motivated by the aforementioned theorems and Corollary 4.4. We also provide numerical experiments, in Section 6, to test these methods.
2. Earlier work
Figure 1. Graph Laplacian $L_{n,t}$ acting on a linear function $f$. Purple color showing positive, and green color negative values of $L_{n,t}f$, where lack of color indicates values near 0
The framework of assuming an underlying low-dimensional manifold of data, in conjunction with graph-related tools and, in particular, the graph Laplacian, has been used extensively. Some examples include work in clustering [3, 4, 5, 6, 7], dimensionality reduction [8, 9], and semi-supervised learning [10].
Several of the approaches to study datasets that use the graph Laplacian leverage that if the manifold is smooth enough and well-behaved, then the graph Laplacian approximates some well-understood operator (for instance, the Laplace-Beltrami operator [11]), which has useful mathematical properties.
Therefore, convergence properties of the graph Laplacian becomes important, and it has been studied in [1, 12, 13, 9]. In particular, and highly influential of this paper, is what the asymptotic convergence looks like near singularities of the manifold, which was shown in [2].
The special case that $\Omega$ is composed of linear manifolds has received earlier attention. For instance, trying to recover facts about the underlying geometry has been explored in [14, 15]. However, these methods have not utilized the graph Laplacian, but other means.
2.1. Main motivating application
There is a lot of interest in understanding the geometry of the loss landscape of neural networks. In particular, zero sets of the loss function.
As a motivating example for the methodology we develop in this paper, consider the following problem. Consider the following class of neural networks:
$f_{W}(x)=\sum_{i=1}^{k}a_{i}(w_{i}\cdot x)_{+},\quad x\in\mathbb{S}^{1},\,a_{i}=\pm 1,\,w_{i}\in\mathbb{R}^{2}$ (2.1)
where $\mathbb{S}^{1}$ is the unit circle. For general spheres this is what is called a spherical neural network and has in essence all the same properties as a standard single hidden layer network. This class of networks has been studied in for instance [16, 17, 18]. Let a target function $g(x)=f_{W^{\ast}}(x)$ be given, but with unknown parameter $W^{\ast}$. We are interested in the set of parameters $\Omega_{\delta}$ such that
$\Omega_{\delta}:=\{W\in\mathbb{R}^{2k}:|f_{W}(x)-g(x)|<\delta,\quad x\in\mathbb{S}^{1}\}.$
We know by the definition of such networks that $\Omega_{\delta}$ is the sublevelset of a piecewise linear function. It is thus clear that the sublevelset is itself a union of linear manifolds.
Declarations of interest
None.
3. Basic mathematical objects and theory
In this section, we provide more precise definitions and introduce the basic mathematical theory we will be using to present and prove our results.
3.1. Conditions on manifolds
We will consider sets of the form $\Omega=\cup_{i}^{m}\Omega_{i}$, where each $\Omega_{i}$ is a smooth and compact $d$-dimensional Riemannian submanifold of $\mathbb{R}^{N}$. We will assume that if any pair $\Omega_{i},\Omega_{j}$, and $i\neq j$, have a non-empty intersection, then this intersection will have dimension lower than $d$.
Our analysis will assume that we only have access to samples from $\Omega$, and so associated to $\Omega$ there is a probability measure with density $p:\Omega\to\mathbb{R}$. For clarity of presentation of our results, we assume that the restriction of $p$ to $\Omega_{i}$ is a uniform distribution. Relaxing this would not significantly alter the main arguments, but the analysis, however, would become more cumbersome.
For any point $x\in\Omega_{i}$ in the interior, we can consider the tangent space $T_{\Omega_{i},x}\simeq\mathbb{R}^{d}$, which we will identify as a subspace of the ambient space $\mathbb{R}^{N}$. More precisely, given open subsets $U\subset\mathbb{R}^{d}$ and $W\subset\Omega_{i}$ ($W$ is open in the subspace topology of $\Omega_{i}$), and a coordinate chart $\alpha:U\to W$ such that $\alpha(0)=x$, we define $T_{\Omega_{i},x}$ as the image of $\mathbb{R}^{d}$ under the action of the Jacobian. We denote the Jacobian $D\alpha:U\to\mathbb{R}^{N\times d}$, evaluated at 0, by $D\alpha(0)$. The best linear approximation to $u\mapsto\alpha(u)$ is, of course, given by $u\mapsto x+D\alpha(0)u$, and $x+T_{\Omega_{i},x}$ is the best flat approximation to $\Omega_{i}$ around $x$.
The definition of $\Omega$ implies that an interior point $x\in\Omega$ can have more than one associated tangent space. For example, if $x\in\Omega_{i}\cap\Omega_{j}$ and $i\neq j$, then both $T_{\Omega_{i},x}$ and $T_{\Omega_{j},x}$ exist, and can be different.
A note on notation is that we denote the interior of a manifold $\Omega_{i}$ by $\operatorname{Int}\Omega_{i}$, and the boundary by $\partial\Omega_{i}$.
3.2. Types of singularities
The following are what we will refer to as singular points, which will be of four different kinds. Given $x\in\Omega=\cup\Omega_{i}$, we have the following types:
  1. (Type 1)
    There is a submanifold $\Omega_{i}$ such that $x\in\partial\Omega_{i}$.
  2. (Type 2)
    There are submanifolds $\Omega_{i}\neq\Omega_{j}$ such that $x\in\operatorname{Int}\Omega_{i}\cap\operatorname{Int}\Omega_{j}$.
  3. (Type 3)
    There are submanifolds $\Omega_{i}\neq\Omega_{j}$ such that $x\in\partial\Omega_{i}\cap\operatorname{Int}\Omega_{j}$.
  4. (Type 4)
    There are submanifolds $\Omega_{i}\neq\Omega_{j}$ such that $x\in\partial\Omega_{i}\cap\partial\Omega_{j}$.
The different types above can of course have non-empty intersection with each other, and a non-singular point is simply a point $x\in\operatorname{Int}\Omega_{i}$ such that if $j\neq i$, then $x\neq\Omega_{j}$. See Fig. 2 for two examples of singularities.
$\Omega_{1}$$\Omega_{2}$
$\Omega_{2}$$\Omega_{1}$
Figure 2. There is a singularity in the intersection of the lines above. The left figure shows a point of Type 4, and the right figure shows a point of Type 2.
3.2.1. $(L,r)$-regular manifolds
To formulate our results for general non-flat manifolds, we will need some measure of how regular, with regard to curvature, our set $\Omega$ is. The following definition captures the necessary information.
Definition 3.1.
Let $\Omega=\cup\Omega_{i}$ be a union of compact submanifolds in $\mathbb{R}^{N}$. We also let $r>0$ be the largest radius such that any point $x\in\operatorname{Int}\Omega$, such that $B_{r}(x)\cap\partial\Omega=\emptyset$, allows coordinate charts $\alpha:U\to B_{r}(x)\cap\Omega_{i}$, where $U\subset\mathbb{R}^{d}$ and $B_{r}(x)\subset\mathbb{R}^{N}$ is an open ball of radius $r$ around $x$. Further, assume also that conditions (A.3) and (A.4) hold over all tuples $(U,W,x,y,\pi,\Omega_{i})$ for some $L>0$. Then we say that $\Omega$ is $(L,r)$-regular.
Example 3.2.
Any smooth and compact submanifold is $(L,r)$-regular for some $r>0$. For instance the graph of the function $x\to x^{2}$ over the compact interval $[-1,1]$ is $(1,1)$-regular.
3.3. Graph Laplacian
In this section we introduce the graph Laplacian and how it acts on real-valued functions defined on $\mathbb{R}^{N}$.
Given $n$ i.i.d. random samples $X=\{X_{1},\ldots,X_{n}\}$ from the distribution with density $p$ on $\Omega$, we build a weighted fully connected graph $G=(V,E)$ as follows: We let each sample $X_{i}$ represent a vertex $i$, and for vertices $i,j\in V$ the weight on $(i,j)\in E$ is given by
$W_{n,t}(i,j):=W_{n,t}(X_{i},X_{j})=\frac{1}{n}K_{t}(X_{i},X_{j})=\frac{1}{n}e^{-\frac{\|X_{i}-X_{j}\|^{2}}{t}}.$
The function $W_{n,t}$ is naturally viewed as an $n\times n$ matrix, and the variable $t$ is in the literature often referred to as the bandwidth of the kernel $K_{t}$.
Remark 3.3.
In the limit analysis as $n\to\infty$, it is useful also normalize by $\frac{1}{t^{d/2+1/2}}$. But, since a priori we do not know the dimension $d$, we will work without this normalization.
We define the diagonal weighted degree matrix as
$D_{n,t}(i,i)=\sum_{j}W_{n,t}(i,j),$
and the graph Laplacian $L_{n,t}$ as
$L_{n,t}=D_{n,t}-W_{n,t}.$
Remark 3.4.
This is often referred to as the unnormalized graph Laplacian. There are other normalizations of this matrix which are used, for example, in [6, 7, 13]. One difference between these normalizations are their limit properties.
Given the fully connected graph $G=(V,E)$, the graph Laplacian above can be seen as an operator acting on arbitrary functions $f:V\to\mathbb{R}$ in the following way:
$L_{n,t}f(X_{i})=\frac{1}{n}\sum_{j}K_{t}(X_{i},X_{j})(f(X_{i})-f(X_{j})),\quad(X_{i},X_{j})\in E.$
We extend this operator to acting on functions that live in the ambient space $f:\mathbb{R}^{N}\to\mathbb{R}$ by the canonical choice
$L_{n,t}f(x)=\frac{1}{n}\sum_{j}K_{t}(x,X_{j})(f(x)-f(X_{j})),\quad x\in\mathbb{R}^{N}.$ (3.1)
Several of our results will be stated in terms of the expected operator:
$L_{t}f(x)=\mathbb{E}_{p}[L_{n,t}f(x)]=\int_{\Omega}K_{t}(x,y)(f(x)-f(y))p(y)\,{\rm d}y.$ (3.2)
Further, an immediate consequence of the linearity of the integral is that
$\begin{split}L_{t}f(x)&=\int_{\Omega}K_{t}(x,y)(f(x)-f(y))\,{\rm d}y\\ &=\sum_{i}\int_{\Omega_{i}}K_{t}(x,y)(f(x)-f(y))p(y)\,{\rm d}y.\end{split}$ (3.3)
This is useful to us, and in our approach we work with the restricted Laplacian $L_{t}^{i}$, which we define as
$L_{t}^{i}f(x)=\int_{\Omega_{i}}K_{t}(x,y)(f(x)-f(y))p(y)\,{\rm d}y.$ (3.4)
3.4. The class of functions
As mentioned in the introduction, we apply the graph Laplacian to linear functions of the form $f(x)=v\cdot x$, where $v$ is a unit vector, and all our results are stated with this assumption.
3.5. A concentration inequality
As stated earlier, many of our results pertain to the expected operator $L_{t}$. To connect these results to the operator $L_{n,t}$, we prove the following concentration inequality:
Theorem 1.
Let $X_{1},\ldots,X_{n}$ be i.i.d. samples from a density $p$ on the union of manifolds $\Omega=\cup\Omega_{i}$. Let $f(x)=v\cdot x$ for $x\in\Omega$ and $v$ is a unit vector. Then
$\mathbb{P}\left(\max_{i}\left|L_{n,t}f(X_{i})-\frac{n-1}{n}L_{t}f(X_{i})\right|>\epsilon\right)\leq 2n\exp\left(-\frac{4e(n-1)\epsilon^{2}}{t}\right).$
This concentration can be improved if we have some further information about the manifold. For instance, if we know that the manifold is flat, and we knew the dimension. This allows us to get tight control over the variance of the graph Laplacian, which is what improves the concentration, but we do not explore this further.
4. Main results
Now that we have the necessary definitions and mathematical background, we are ready to present our main results. Before stating the theorems, we will provide a brief section that explains the geometry of some terms that will be used.
General structure of results
By 3.3 it is enough to understand the restricted Laplacian, $L^{i}_{t}$ defined in 3.4. Because of this, our results are formulated to show the behavior of $L^{i}_{t}$. Depending on what type of singularity being examined, it is straight forward to extend the results to the full Laplacian. In Corollary 4.4 we give one example of how to extend the results to the sum $\sum_{i=1}^{2}L^{i}_{t}$ when one is close to an intersection of two manifolds.
Geometry and notation for Section 4.1
We will in several theorems also formulate the function $x\to L_{t}^{i}f(x)$ partly in terms of new coordinates $(r,\theta)$. Here $r$ is defined by the relation $\left\lVert x-x_{0}\right\rVert=\sqrt{t}r$, and given the projection $\hat{x}$ of $x$ to the tangent plane of the flat manifold $\Omega_{i}$, we define $\theta\in[0,\pi/2]$ to be the angle between vectors $x-x_{0}$ and $x_{0}-\hat{x}$, as the schematic in Fig. 3. By simple geometry, it follows that $\left\lVert\hat{x}-x\right\rVert=r\sqrt{t}\sin\theta$.
Given a vector $v\in\mathbb{R}^{N}$, we will have reason to write the expression $v\cdot(\hat{x}-x)$ as
$v\cdot(\hat{x}-x)=r\sqrt{t}\sin(\theta)v\cdot\frac{\hat{x}-x}{\left\lVert\hat{x}-x\right\rVert}=r\sqrt{t}\sin(\theta)v_{n,\Omega_{i}}(x),$
where we have defined
$v_{n,\Omega_{i}}(x)\coloneqq\begin{cases}v\cdot\frac{\hat{x}-x}{|\hat{x}-x|}&\text{if }x\neq\hat{x}\\ 0&\text{if }x=\hat{x}.\end{cases}$
Here $v_{n,\Omega_{i}}$ is the projection of $v$ onto a unit normal vector of $\Omega_{i}$, and which can vary with $x$. We will later use a natural coordinate system parallel to $\Omega_{i}$, see A.1, such that when $x\neq\hat{x}$, this function is constant up to its sign. This implies that evaluating $r\sqrt{t}\sin(\theta)v_{n,\Omega_{i}}(x)$ is the same as letting $v_{n,\Omega_{i}}$ be fixed, but allowing $\theta$ to change sign depending on which side of $\Omega_{i}$ $x$ is, i.e. as if we have fixed the coordinate system in which we measure the angle $\theta$. In the following, to increase readability, we suppress the $x$-dependency of $v_{n,\Omega_{i}}$
Additionally, in Theorem 3 we will have a term $v_{n,\partial\Omega_{i}}$ that is specific to that theorem. This will be defined in the case where there is a boundary close to $x$. In Fig. 3, this would imply there is a boundary of $\Omega_{1}$ nearby. To give the definition of this term first assume that $\Omega_{i}$ is a flat manifold and assume that close to the point $x$, $\partial\Omega_{i}$ is flat. Then, we first let $\hat{x}_{\partial\Omega_{i}}$ be the projection of $\hat{x}$ to $\partial\Omega_{i}$. We can now define a unit normal at $\hat{x}_{\partial\Omega_{i}}$, denoted by $n_{\partial\Omega_{i}}$. Two choices are natural, a normal pointing either towards, or away from $\Omega_{i}$. We define $n_{\partial\Omega_{i}}$ as the latter. Given a vector $v\in\mathbb{R}^{N}$, we define
$v_{n,\partial\Omega}\coloneqq v\cdot n_{\partial\Omega}.$
Since we assume that close to $x$, $\partial\Omega$ is flat, $n_{\partial\Omega}$ can be considered constant.
$x_{0}$$x$$\sin\theta r\sqrt{t}$$\hat{x}$$\theta$$r\sqrt{t}$$\Omega_{1}$$\Omega_{2}$$r_{0}\sqrt{t}$
Figure 3. Schematic picture of the geometry of Theorem 2, where $\Omega_{1}$ is the object of interest and $x\in\Omega_{2}$ for visualization purposes.
Geometry and notation for Section 4.2
To help with the geometric picture for general manifolds, the situation is as explained in Section 4 and Fig. 3: the terms $x,x_{0},\hat{x},\theta$ and $v_{n,\Omega_{i}}$ are in the same relation to each other as in Section 4. The main difference is that the tangent space now changes with $x$ as the manifold is no longer flat. In that sense the geometry for more general manifolds is not more difficult, but handling error terms is more involved.
4.1. Flat manifolds
In this section we assume that $\Omega=\cup_{i}\Omega_{i}$, where each $\Omega_{i}$ is a flat manifold. This means that each coordinate chart around a point $x\in\operatorname{Int}\Omega_{i}$ is an isometry with an open neighborhood $U$ of $x$, where $U$ is a ball in $\mathbb{R}^{d}$.
In Theorem 2 we give a result concerning the behavior of $x\to L_{t}^{i}f(x)$ when we are not close to the boundary $\partial\Omega$. This case is easier to prove, and we give explicit bounds of all terms involved, and express them with elementary functions.
In Theorem 3 we show what happens when we are close to $\partial\Omega$, and here our results contain more complex expressions for some terms.
In the following theorems, we think of $x_{0}$ as the potentially singular point, see Fig. 3. The appearance of terms different from zero either comes from the fact that we are evaluating the restricted operators outside of their corresponding manifolds, or because that we are close to the boundary. For example, if $\Omega=\Omega_{1}\cup\Omega_{2}$ and $x_{0}\in\Omega_{1}\cap\Omega_{2}$, the extra contribution thus comes from each $L_{t}^{i}$, $i\in\{1,2\}$, when evaluated on the other manifolds, i.e. $L_{t}^{1}f(x)$ when $x\in\Omega_{2}$ and vice versa. By combining Theorem 2 and Theorem 3, it is possible to consider several types of singularities defined in Section 3.2.
Theorem 2.
Let $f(x)=v\cdot x$ for some unit vector $v\in\mathbb{R}^{N}$ and assume that $p$ is the uniform density over $\Omega_{i}$. Let $x_{0}\in\Omega_{i}$ and assume that $\partial\Omega_{i}\cap B_{2R}(x_{0})=\emptyset$ for $R=r_{0}\sqrt{t}$, where $r_{0}>2$. Further, $x\in B_{R}(x_{0})$, and $v_{n,\Omega_{i}}$, $r$ and $\theta$ are as described in Section 4. If $t\leq\frac{R^{2}}{d/2+1}$, $d\geq 1$ and $r<r_{0}/2$, then we have that
$L^{i}_{t}f(x)=t^{d/2+1/2}\left(A(\theta,r_{0},d)v_{n,\Omega_{i}}\sin\theta re^{-\sin^{2}\theta r^{2}}+B(x)e^{-r_{0}^{2}}\right),$
where $A,B$ are real-valued functions. The function $B$ depends on $x$, but it is uniformly bounded by $|B(x)|\leq\frac{d+1}{4}r_{0}^{d}\frac{|\mathbb{S}^{d-1}|}{|\Omega_{i}|}$; and $A$ depends on $x$ only through $\theta$, and is bounded by
$\frac{1}{2}\max(\pi^{d/2},2\pi^{d/2}-|\mathbb{S}^{d-1}|2^{d/2}r_{0}^{d-1}e^{-r_{0}^{2}+1})\leq A(\theta,r_{0},d)\leq\pi^{d/2}.$
The following theorem is an extension of Theorem 2 to the case when the ball $B_{R}(x_{0})\cap\partial\Omega_{i}\neq\emptyset$, which gives rise to an additional term in the expression of $L_{t}^{i}f(x)$. We again refer to the schematic picture of Fig. 3 and comments in Section 4 for explanation of the coordinates $(r,\theta)$, function $v_{n,\Omega_{i}}$ and constant $v_{n,\partial\Omega_{i}}$.
Theorem 3.
Let $f(x)=v\cdot x$ for some unit vector $v\in\mathbb{R}^{N}$, and assume that $p$ is the uniform density over $\Omega_{i}$. Let $x_{0}\in\Omega_{i}$ and assume that $\partial\Omega_{i}\cap B_{2R}(x_{0})$ is part of a $d-1$ dimensional plane for $R=r_{0}\sqrt{t}$, where $r_{0}>2$. Further, $x\in B_{R}(x_{0})$, and $v_{n,\Omega_{i}}$, $v_{n,\partial\Omega_{i}}$, $r$ and $\theta$ are as described in Section 4. If $t\leq\frac{R^{2}}{d/2+1}$, $d\geq 1$ and $r<r_{0}/2$, then we have that
$L^{i}_{t}f(x)$ $=\widehat{A}_{1}(x)t^{\frac{d+1}{2}}v_{n,\Omega_{i}}\sin(\theta)re^{-\sin^{2}(\theta)r^{2}}+\widehat{A}_{2}(x)t^{\frac{d}{2}}v_{n,\partial\Omega_{i}}e^{-\sin^{2}(\theta)r^{2}}$
$\quad+B(x)t^{\frac{d+1}{2}}e^{-r_{0}^{2}},$
for explicitly computable function $\widehat{A}_{2}$, and with explicitly computable bounds of function $\widehat{A}_{1}$. The function $B$ has the same bounds as in Theorem 2.
Remark 4.1.
The function $\widehat{A}_{1}$ is bounded by
$\frac{1}{2\delta_{0}}\left(e^{-k_{0}^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-k_{0}^{2}\right)-\frac{2(\delta_{0}^{2}-k_{0}^{2})^{\frac{d-1}{2}}}{d-1}\right)\leq\widehat{A}_{1}\leq\Gamma\left(\frac{d-1}{2}\right)\sqrt{\pi}$
and $\widehat{A}_{2}$ is given by
$\widehat{A}_{2}=\frac{|\mathbb{S}^{d-2}|}{2}\left(e^{-\delta_{0}^{2}}\frac{(\delta_{0}^{2}-k_{0}^{2})^{(d-1)/2}}{d-1}+\frac{1}{2}e^{-k_{0}^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-k_{0}^{2}\right).\right)$
To define $k_{0}$ and $\delta_{0}$, we recall the geometric picture of Section 4. Then $K$ is the projection of $(\hat{x}-\hat{x}_{\partial\Omega_{i}})$ to $n_{\partial\Omega_{i}}$, $k_{0}=K/\sqrt{t}$, and $\delta_{0}=\sqrt{r_{0}^{2}-r^{2}\sin^{2}\theta}$.
4.2. General manifolds
In this section we will not assume the manifolds are flat, instead we assume that $\Omega$ is $(L,2R)$-regular, see Definition 3.1. The type of singularity we deal with for a more general manifold will be a Type 2, and we will assume we are not too close to any boundary. Although we achieve a similar result as in Theorem 2, the constants depend on the curvature of the manifold which in many practical applications would be unknown. This comes from the non-local nature of the $t$-regularized Laplacian $L_{t}$. It is however useful if we are in a situation where some a-priori information about the curvature is known, for instance if we are interested in exploring level sets of certain functions.
Theorem 4 (General manifold).
Let $f(x)=v\cdot x$ for some unit vector $v\in\mathbb{R}^{N}$ and assume that $p$ is the uniform density over $\Omega_{i}$, a $(L,2R)$-regular manifold. Let $x_{0}\in\Omega_{i}$ and assume that $\partial\Omega_{i}\cap B_{2R}(x_{0})=\emptyset$ for $R=r_{0}\sqrt{t}$, where $r_{0}>2$. Further, $x\in B_{R}(x_{0})$, and $v_{n,\Omega_{i}}$, $r$ and $\theta$ are as described in Section 4. If $L4R^{2}/\sqrt{t}\leq\frac{1}{2}$, $t\leq\frac{R^{2}}{d/2+1}$, $d\geq 1$ and $r<r_{0}/2$, then we have that
$\begin{split}L_{t}^{i}f(x)=t^{d/2+1/2}\widehat{A}(x)v_{n,\Omega_{i}}r\sin\theta e^{-r^{2}\sin^{2}\theta}+t^{d/2}C_{L,R}(x)4p\pi^{d/2}+D(x)e^{-r_{0}^{2}}.\end{split}$
In the above, $\widehat{A}>0$ is a function such that
$|A(d,r,\theta)-\widehat{A}(x)|\leq(1+3C_{L,R})A(d,r,\theta)$
where $A(d,r,\theta)$ as in Theorem 2; $C_{L,R}$ is a function such that
$|C_{L,R}(x)|\leq 4(L4R^{2})^{2}R/t+LR^{2}(1+4LR^{2})$
and $|D(x)|\leq\operatorname{diam}(\Omega)$.
The next lemma gives useful bounds on $L^{i}_{t}f(x)$ when $x$ is non-singular.
Lemma 4.2.
Given the conditions of Theorem 4 and the additional assumption that $x\in\Omega_{i}$, we have that
$L_{t}^{i}f(x)=t^{d/2}\widehat{A}(x)LR^{2}+t^{d/2}C_{L,R}(x)4p\pi^{d/2}+D(x)e^{-r_{0}^{2}},$
Remark 4.3.
The result in Lemma 4.2 can be used together with both Theorem 2 and Theorem 4 to analyze the behavior of the mapping $x\to L_{t}f(x)$ around intersections.
The following corollary is a consequence of Theorems 4 and 4.2. The geometry is as in Section 4, projecting $x$ specifically to the tangent plane $T_{\Omega_{1},x_{0}}$.
Corollary 4.4.
Let $f(x)=v\cdot x$ for some vector $x\in\mathbb{R}^{N}$ and assume that $p$ is the uniform density over a $(L,2R)$-regular manifold $\Omega=\cup_{i=1}^{2}\Omega_{i}$. Let $x_{0}\in\Omega_{1}\cap\Omega_{2}$ and assume that $\partial\Omega_{i}\cap B_{2R}(x_{0})=\emptyset$ for $i\in\{1,2\}$ and $R=r_{0}\sqrt{t}$, where $r_{0}>2$. Assuming $R$ and $t$ satisfies the conditions of Theorem 4, then for $x\in B_{R}(x_{0})\cap\Omega_{2}$ such that $\|x-x_{0}\|=r\sqrt{t}$ for $r<r_{0}/2$, we have that
$\begin{split}L_{t}f(x)&=t^{d/2+1/2}\widehat{A}(x)v_{n,\Omega_{1}}r\sin\theta_{1}e^{-r^{2}\sin^{2}\theta_{1}}+t^{d/2}\widehat{A}(x)LR^{2}\\ &\qquad+t^{d/2}C_{L,R}(x)8p\pi^{d/2}+2e^{-r_{0}^{2}}D(x)\end{split}$
In the above, $\theta$ and $v_{n,\Omega_{1}}$ are as in Section 4, with $\Omega_{i}=\Omega_{1}$. Functions $\widehat{A},C_{L,R}$ and $D$ are as in Theorem 4.
4.3. Manifolds with noise
In earlier sections, we have assumed that the samples used to evaluate $L_{n,t}f(x)$ are taken directly from $\Omega$. However, in many real-world applications we can only expect that we access the points $X_{1},X_{2},\dots X_{n}$ approximately, for instance due to measurements uncertainties. In this section we propose a scenario where our results can have application also in this more general case. For instance, we could have the following situation:
  1. (1)
    We first obtain a set observations $x_{1},x_{2},\dots x_{n}$ of $X_{1},X_{2},\dots,X_{n}$ from $\Omega$.
  2. (2)
    We cannot observe these points directly, but instead we are able to take multiple, noisy, measurements from the random variables $x_{1}+\epsilon_{1},x_{2}+\epsilon_{2},\dots x_{n}+\epsilon_{n}$, where each $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}I)$
To model this is, we replace the operator
$L_{n,t}f(x)=\frac{1}{n}\sum_{j=1}^{n}K_{t}(x,X_{j})(f(x)-f(X_{j})),$
by
$L_{n,t,\epsilon}f(x)=\frac{1}{n}\sum_{j=1}^{n}K_{t}(x,X_{j}+\epsilon_{j})(f(x)-f(X_{j}+\epsilon_{j})).$
The following theorem shows that with our results, and with regards to the second source randomness, the uncertainty in measurements, we can understand the expectation of $L_{n,t,\epsilon}$.
Theorem 5 (Stochastic version).
Let $L_{n,t,\epsilon}$ be as above, and the operator $\mathbb{E}_{\epsilon}[\,\cdot\,]=\mathbb{E}[\,\cdot\mid X_{1},\dots,X_{N}]$ be expectation with regard to the random variables $(\epsilon_{1},\dots,\epsilon_{n})$. Then
$\mathbb{E}_{\epsilon}L_{n,t,\epsilon}f(x)=\frac{2t^{N/2+1}}{(2\sigma^{2}+t)^{N/2+1}}\frac{1}{n}\sum_{j=1}^{n}K_{2\sigma^{2}+t}\left(x,X_{j}\right)(f(x)-f(X_{j})).$
Thus, for $t^{\prime}=t+\sigma^{2}$ and up to normalization, $\mathbb{E}_{\epsilon}L_{n,t,\epsilon}$ is the opeartor as $L_{n,t^{\prime}}$.
5. Hypothesis test for singularities in flat manifolds
In this section we will develop a hypothesis test for singularities in flat manifolds. We begin by assuming that $\Omega=\cup_{i}\Omega_{i}$ is a union of flat manifolds, and that we have a set of samples $\{X_{1},\ldots,X_{n}\}$ from the uniform density $p$ on $\Omega$. In this section we wish to consider the following hypothesis for a fixed point $x_{0}\in\Omega$:
  • $H_{0}$: $B_{2R}(x_{0})\cap\Omega=B_{2R}(x_{0})\cap\Omega_{i}$ for some $i$ and $B_{2R}(x_{0})\cap\partial\Omega_{i}=\emptyset$.
  • $H_{1}$: Hypothesis $H_{0}$ is false.
Theorem 6.
Let $\Omega=\cup\Omega_{i}\subset\mathbb{R}^{N}$ be a union of flat manifolds and let $f(x)=v\cdot x$ for some unit vector $v\in\mathbb{R}^{N}$. Assume that $p$ is the uniform density over $\Omega$. Let $X_{1},\dots,X_{n}$ i.i.d. samples from $p$, define the statistic
$T=\max_{m:X_{m}\in B_{1}(x_{0})}|L_{n,t}f(X_{m})|,$
and for a fixed test-level $\alpha\in(0,1)$ the rejection region $\Theta:=\{T>\delta\}$, where
$\delta=\sqrt{\frac{t}{e(n-1)}\log\left(\frac{2n}{\alpha}\right)}.$
If
$t\leq\min\left\{1,\frac{2}{\log\left(\frac{e(n-1)}{2\log\left(\frac{2n}{\alpha}\right)}\right)}\right\},$
and we reject $H_{0}$ in the rejection region $\Theta$, then this is a test of level $\alpha$.
Remark 5.1.
Under the null hypothesis that $\Omega_{i}$ is $(L,R)$-regular, we can use the same test statistic $T$, but we change our rejection region by redefining $\delta$ using Theorem 4. This $\delta$ will necessarily be larger and thus if we wish to detect singularities in more general manifolds, we need more samples to achieve the same power.
As with all hypothesis tests we can only safely reject the null hypothesis, but why we reject it is not clear. Looking at Theorem 3 it could either be that we have a singularity of the types Type 1Type 4 if we assume that all submanifolds are flat. However, the reason for rejection could also be that the manifold is not flat. What we will outline next is a way to estimate the power of the test in Theorem 6 in the case when the alternative hypothesis is that there is an intersection between two flat manifolds. Similar results can be obtained for a singularity of any of the types Type 1Type 4.
Theorem 7.
Let $\Omega=\Omega_{1}\cup\Omega_{2}$ and otherwise as in Theorem 6. Let $H_{1}^{\prime}$ be an alternative hypothesis that there is an intersection at $x_{0}$ between $\Omega_{1}$ and $\Omega_{2}$ where both are flat manifolds of intrinsic dimension $d$, where $x_{0}$ is a singular point of Type 2 and $\partial\Omega_{i}\cap B_{2R}(x_{0})=\emptyset$, $i=1,2$. Let $T$ be the test statistic in Theorem 6 and $\delta$ be as in Theorem 6. Then the probability of rejecting $H_{0}$ under $H_{1}^{\prime}$ (power) is bounded by
$\mathbb{P}(T>\delta\mid H_{1}^{\prime})\geq 1-\alpha-\mathbb{P}(X_{m}\in B^{c}_{R/3}(x_{0}))^{n},$
when
$t(n,\alpha)=\min\left\{1,\frac{2}{\log\left(\frac{e(n-1)}{\log\left(\frac{2n}{\alpha}\right)}\right)}\right\},$
and $n$ is so large that
$\begin{split}t\left((d+1)\log(1/t)+\log\left(\frac{2}{\pi^{d}|v_{n,\Omega_{1}}|^{2}}\right)\right)&\leq 2(1-\sin^{2}\theta_{1}),\\ t&\leq\frac{1}{d/2+\log\left(\frac{16^{2}e^{2}}{\pi^{d}|v_{n,\Omega_{1}}|^{2}\sin^{2}\theta_{1}}\right)}.\end{split}$ (5.1)
Remark 5.2.
Asymptotically as $|v_{n,\Omega_{1}}|$ and $\sin\theta$ tends to zero, the number of samples needed to satisfy 5.1 is
$\frac{n}{\log n}=O\left(\frac{1}{|v_{n,\Omega_{1}}|^{4}\sin^{4}\theta_{1}}\right).$
As such we see that we drastically loose power if the angle $\theta_{1}$ is small, or if the normal vector $v_{n,\Omega_{1}}$ is small. This is intuitive, as the smaller $|v_{n,\Omega_{1}}|$ is, the less we see the intersection in the chosen direction $v$ and if it is zero, the Laplacian applied to $f$ will not see the intersection at all. On the other hand, if $\theta_{1}$ is small, the intersection becomes less distinct. When $\theta=0$, the two manifolds are tangent, and from the perspective of the Laplacian (for small $t$), they are effectively the same manifold, as such there is no singularity. In a practical application we can search for a vector $v$ such that $T$ is maximized (on a subsample). Thus, the only real constraint is the actual angle of intersection between the manifolds.
6. Numerical Experiments
6.1. Hypothesis test for singularities
In this section we will demonstrate the hypothesis test in Theorem 6 for a flat manifold. We will assume that we have a set of samples $(X_{1},X_{2},\ldots,X_{n})$ distributed according to the uniform density on $\Omega$, and that we have a fixed point $x_{0}\in\Omega$. We will assume that the null hypothesis $H_{0}$ is true, and that $x_{0}$ is not close to any boundary. Specifically we will assume that $N=3$ and that $\Omega$ is piece of a hyperplane in $\mathbb{R}^{3}$. We will then evaluate $L_{n,t}f(x)$ on $(X_{1},X_{2},\ldots,X_{n})$ and calculate the test statistic $T$ as in Theorem 6, with $t$ as in Theorem 7. We will then calculate the rejection region $\Theta$ and see if $T$ is in $\Theta$. You will find the results in Fig. 4.
To test the power of Theorem 7 we construct an intersection between two hyperplanes in $\mathbb{R}^{3}$ at different angles. The result is in Fig. 4.
Samples Rejections under $H_{0}$ Rejections under $H_{1}^{\prime}$ ($\theta=\pi/4$) $H_{1}^{\prime}$ ($\theta=\pi/2$)
$2\cdot 10^{4}$ 00 00 $0.04$
$3\cdot 10^{4}$ 00 00 $0.73$
$4\cdot 10^{4}$ 00 00 $1$
$5.5\cdot 10^{4}$ 00 $0.2$ $1$
$6\cdot 10^{4}$ 00 $0.28$ $1$
$6.5\cdot 10^{4}$ 00 $0.57$ $1$
$7\cdot 10^{4}$ 00 $0.81$ $1$
Figure 4. Results of hypothesis test for singularities in flat manifolds, each experiment run 100 times.
6.1.1. Neural networks
As mentioned in the introduction (see Section 2.1) we are interested in the zero set of a neural network. Specifically, let us consider the zero set of a neural network as in Section 2.1, where $k=3$ and the target function $f_{W^{\ast}}$ is chosen such that all three weights $w_{i}$ are the same and $a_{1}=-a_{2}=-a_{3}$ and $\delta=0.01$. This ensures that the zero set is large as due to the cancellations in the sum there should be many weights that produce the same function $f_{W^{\ast}}$. For details how the dataset was constructed see Appendix F. In Fig. 5 we see the zero set of the neural network presented using PCA from 6 to 3 dimensions. This suggests that the zero set is composed of the intersection of piecewise flat manifolds, and we can see the result of applying the graph Laplacian to a linear function ($v$ that was chosen to maximize the operator). The direction was chosen on an independent subset and the test performed on another independent subset of $30000$ samples.
Figure 5. Zero set of a neural network in $\mathbb{R}^{6}$ projected to $\mathbb{R}^{3}$ using PCA on the left and the graph Laplacian applied to a linear function on the right clearly showing the functional form $xe^{-x^{2}}$. The dashed line corresponds to the rejection region $\Theta$ in Theorem 6 and $t$ is chosen according to Theorem 7.
6.2. Estimating singularities
In the following experiments we demonstrate how to estimate the location of an intersection and the intersecting angle $\theta$ of a union of two manifolds, given that we know that $\Omega=\Omega_{1}\cup\Omega_{2}$, and $\Omega_{1}\cap\Omega_{2}$ is non-empty and we have a Type 2 singularity. Further, we assume that we have both a set of samples $X\subset\Omega$ distributed according to the associated density on $\Omega$, and an additional set of points $Y$ from a curve $\Gamma\subset\Omega_{i}$, for some $i\in\{1,2\}$. The curve $\Gamma$ intersects $\Omega_{1}\cap\Omega_{2}$, and we assume that no other singularity is very close, which is a situation like in Fig. 7 and Fig. 10.
This particular setup could, for instance, be understood in the context of urban traffic flow analysis, where $\Omega_{1}$ and $\Omega_{2}$ represent two intersecting roads in a city, or some kind of interchange. In this scenario,
  • The set $X$ represents GPS data points from different vehicles traveling on the roads or interchange.
  • The set $Y\subset\Gamma$ represents samples from the path of a specific vehicle traveling through the intersection or interchange.
  • The intersection $\Omega_{1}\cap\Omega_{2}$ is the point of intersection between the roads, and the properties of $\Omega_{1}\cap\Omega_{2}$ could be of interest for understanding traffic flow.
[…]
Figure 6. Graph of $y=xe^{-x^{2}}$
Now let $Y=\{y_{1},\dots,y_{m}\}\subset\Gamma$ be as above. When we evaluate $x\mapsto L_{n,t}f(x)$ on $Y$, we get the values $P=\{p_{1},\dots,p_{m}\ :\ p_{i}=L_{n,t}f(y_{i}),\ p_{i}\in\Gamma\}$. From our previous results, we know that these values, with enough samples, are close to
$L_{t}f(y_{i})=t^{d/2+1/2}\left(A(d,r_{0},\theta)v_{n,\Omega_{i}}\sin\theta_{i}re^{-\sin^{2}\theta r_{i}^{2}}\right)+\text{Error}(y_{i},r_{0},L).$ (6.1)
The function depends on $r=\left\lVert y_{i}-x_{0}\right\rVert$, and the angle $\theta=\angle(y_{i}-x_{0},\hat{y}_{i}-x_{0})$, where $\hat{y}_{i}$ is the projection of $y_{i}$ onto either, in the flat case, a manifold $\Omega_{i}$ or, in the curved case, its tangent plane, as defined in Theorem 4, for some point $x_{0}\in\Omega_{i}$. The error term depends on $x$,$r_{0}$ and $L$. As before, the constant $L$ bounds the curvature of $\Omega$. The error term can be bounded using the results in Section 4.1 and Section 4.2. By choosing a sufficiently small $t$, we can make $r_{0}$ arbitrarily large, causing the function $A(d,r_{0},\theta)$ to approach $2\pi^{d/2}$. For our purposes in this section, we assume that $A(d,r_{r}0,\theta)\approx 2\pi^{d/2}$ and that $\text{Error}(x,r_{0},L)\approx 0$ are good enough approximations. With these assumptions,
$L_{t}f(y_{i})=t^{d/2+1/2}2\pi^{d/2}v_{n,\Omega_{i}}\sin\theta_{i}re^{-\sin^{2}\theta r_{i}^{2}}.$ (6.2)
Remark 6.1.
Theorem 4 uses $\widehat{A}(x)$ instead of $A(d,r_{0},\theta)$, but $\widehat{A}(x)$ can be made arbitrarily close to $A(d,r_{0},\theta)$ by choosing $t$ or $L$ small enough. How small we can make $L$, however, depends on how well-behaved the manifold is.
We will use the functional form of equation 6.2 and our two sets of samples to get estimators to the location and the angle of intersection of our manifold. We do not give explicit probabilistic bounds here, but assume $L_{n,t}f$ is sufficiently close to $L_{t}$.
Remark 6.2.
Let us note that a similar concentration bound as in Theorem 1 should be possible for this type of scenario, which would allow us to get explicit bounds for these experiments. That and not using the approximations above, would allow us to get explicit probabilistic intervals for all estimators.
6.2.1. Estimators for Intersection Location and Angle
In this section we derive estimators for the location of the intersection and angle of intersection, based on 6.2.
Let $g(r,\theta)=v_{n,\Omega}r\sin\theta e^{-r^{2}\sin^{2}\theta}$. We observe that $g$ only depends on $x=r\sin\theta$ (up to the sign of $v_{n,\Omega_{i}}$), allowing us to write $g(r,\theta)=g(x)$, with $g(x)$ being a rescaled version of the function $h(x)=xe^{-x^{2}}$. See Fig. 6 for an illustration of $h$.
The function $h$ attains its minimum and maximum values at $z_{1}=-\frac{1}{\sqrt{2}}$ and $z_{2}=\frac{1}{\sqrt{2}}$ respectively. The intersection corresponds to the midpoint of these two points. We can estimate the point $s$ where $\Gamma$ intersects $\Omega_{1}\cap\Omega_{2}$ as:
$\hat{s}=\frac{\arg\max_{x_{i}}(P)+\arg\min_{x_{i}}(P)}{2}.$
To estimate the intersection angle $\hat{\theta}$ of $\theta$, we first need $\hat{r}_{\max}$, an estimate of the scaled distance from 0 that maximizes $g(r,\theta)$. This is
$\hat{r}_{\max}=\frac{\left\lVert\hat{s}-\arg\max_{x_{i}}P\right\rVert}{\sqrt{t}}.$
Since $\max_{r}g(r,\theta)=\frac{1}{\sqrt{2}}$, we have that $\hat{r}_{\max}\sin\theta\approx\frac{1}{\sqrt{2}}$. Solving for $\theta$, this leads to our estimator:
$\hat{\theta}=\arcsin\left(\frac{1}{\sqrt{2}\hat{r}_{\max}}\right).$
6.2.2. Experimental setup
We tested our estimation methods on hypersurfaces in $\mathbb{R}^{3}$ intersecting at angles given by $\theta\in\{\pi/2,\pi/4,\pi/8,\pi/16\}$, and with $t=10^{-3}$. For both flat and curved manifolds, we conducted 100 runs with random chosen functions of $f(x)=v\cdot x$, where $v$ is sampled from the uniform distribution on $\mathbb{S}^{2}$.
For each run:
  • We sampled $2\times 10^{4}$ points, from both $\Omega_{1}$ and $\Omega_{2}$, in a bounded region near the intersection
  • We evaluated $L_{n,t}$ on $10^{3}$ uniformly spaced points on $\Gamma$ to obtain our set $P$.
Remark 6.3.
We chose $\mathbb{R}^{3}$ mainly for visualization purposes. The intrinsic dimension of $\Omega$ remains low, regardless of the ambient dimension. However, choosing a function $f(x)=v\cdot x$ for $L_{t}^{i}$ to act on becomes more challenging in higher dimensions, as it becomes harder to orient $v$ such that $v_{n,\Omega_{i}}$ is large, especially when $v$ is chosen randomly, as in our experiments.
In the following sections we present our results for each type of manifold and angle of intersection.
6.2.3. Results for flat manifolds
In this section, we apply our estimation methods to the case that $\Omega_{1},\Omega_{2}$ are flat. Since the manifolds are flat, the angle $\theta_{i}=\theta$, where $\theta\in[0,\frac{\pi}{2}]$ is fixed.
In Fig. 7 we see our samples of $\Omega$ and $\Gamma$, and in Fig. 8 we see an example of the values we get in $P$. Finally, in Fig. 9 we see how well this approach works in trying to learn both $\theta$ and $r$.
Figure 7. Samples of $\Omega=\Omega_{1}\cup\Omega_{2}$ and $\Gamma$ (blue), where $\Omega_{1}$ (green) and $\Omega_{2}$ (red) have no curvature.
Figure 8. $L_{n,t}f$ evaluated on $\Gamma$. Flat manifolds.
Figure 9. Estimates of $\theta$ and $s$ on flat manifolds.
6.2.4. Results for curved manifolds
Here we test these methods in the case that $\Omega=\Omega_{1}\cup\Omega_{2}$ is $(L,2R)$-regular, with $L=0.5$ and $R$ having no upper bound. The setup is the same as what we see in Fig. 10.
If we let $\theta$ denote the angle between the tangent spaces of $\Omega_{1}$ and $\Omega_{2}$ at $x_{0}$, then since there is curvature, we cannot expect $\theta_{i}=\theta$ for every $i=1,\dots,n$, or even between any pair of them. However, we can still estimate the location of the intersection as before, and estimating $\theta$ in this way provides some information about the intersection, even if it is not as strong as in the case without curvature. The range of possible values for $\theta$, due to curvature, can be bounded by knowing the bounds of the curvature constant $L$.
In Fig. 10 we see our samples of $\Omega$ and $\Gamma$, in Fig. 11 we see an example of the values we get in $P$. Finally, in Fig. 12 we see how well this approach works in trying to learn both $\theta$ and $r$.
Figure 10. Samples of $\Omega=\Omega_{1}\cup\Omega_{2}$ and $\Gamma$ (blue), where $\Omega_{1}$ (green) and $\Omega_{2}$ (red) have curvature.
Figure 11. $L_{n,t}f$ evaluated on $\Gamma$. Curved manifolds.
Figure 12. Estimates of $\theta$ and $s$ on curved manifolds.
7. Final remarks
In this paper we built upon the work of [2] and developed explicit versions of their asymptotic analysis of $x\to L_{t}f(x)$. Our results are the strongest and most useful in the case of flat manifolds, and the motivation to focus on this scenario comes partly from neural networks (see Section 2.1).
We used these explicit bounds to construct hypothesis tests with theoretical guarantees on the level of the test as well as the power under specific conditions. The numerical experiments in Section 6.1 suggests that the power can be close to $1$ with many samples as predicted by Theorem 7. The power of these hypothesis tests can clearly be seen in the example of the zero set of a neural network (see Section 6.1.1) where we can safely reject the null hypothesis that the zero set is a single flat manifold.
While the bounds in Theorem 4 are weaker, our numerical experiments suggest that this approach can be useful for gaining geometric information about the union of more general manifolds $\Omega=\cup_{i}\Omega_{i}$. In [2], the authors mainly considered sets $\Omega=\cup_{i=1}^{n}\Omega_{i}$ with $n\leq 2$. Our approach of splitting $L_{t}$ into components $L_{t}^{i}$ makes it straight forward to directly apply our theorems for $n\geq 2$, allowing us to consider a wider range of singularities. For example, we can extend the framework to examine points that are both of Type 1 and Type 2, or to study intersections of more than two manifolds.
In our numerical experiments in 6.2, we assumed that $\Omega=\cup_{i=1}^{2}\Omega_{i}$ and had access to samples of an intersecting curve $\Gamma$, which allowed us to estimate geometric properties near intersections. Future work could involve extending our framework to other types of singularities and developing similar tests and estimators. It would also be interesting to explore methods that do not rely on access to such curves.
Similar theorems can be proven for other kernels besides the Gaussian one, as many ideas used in our proofs are not specific to the Gaussian case, but rather rely mainly on symmetries of $K_{t}$. Investigating the use of other kernels and comparing their performance in different scenarios is a potential direction for future research.
Acknowledgments
The first author was supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The second author was supported by the Swedish Research Council grant dnr: 2019–04098.
Appendix A Preliminaries on manifolds and explicit bounds on Gamma functions
A.1. Integration on $\Omega$
We will integrate scalar-valued functions, $f:\Omega\to\mathbb{R}$, over $\Omega$. When formulating integration of scalar-valued functions over submanifolds of $\mathbb{R}^{N}$, we follow the approach in [19]. Because we need some preliminary results concerning integration on $\Omega$, we make some important definitions explicit.
First, let $x_{1},\ldots,x_{k}$ be vectors in $\mathbb{R}^{N}$ for $k\leq N$. If $I=(i_{1},i_{2},\dots,i_{k})$ is a $k$-tuple of integers such that $i_{1}\leq i_{2}\leq\cdots\leq i_{k}$, define $X_{I}\in\mathbb{R}^{k\times k}$ as the $k\times k$ matrix containing only rows $i_{1},\dots,i_{k}$ of the matrix $X=(x_{1},\ldots,x_{k})$. Now we can define the volume function $V:\mathbb{R}^{N\times k}\to\mathbb{R}$, by $V(X)=\sqrt{\det^{2}(X^{t}X)}=\left[\sum_{I}\det^{2}X_{I}\right]^{1/2}$, where the $I$’s span over $k$-tuples as above, see [19, Theorem 21.4].
In general, given a coordinate chart $\alpha:U\to W$, where $U\subset\mathbb{R}^{d}$, $W\subset\Omega_{i}\subset\mathbb{R}^{N}$ are open subsets, and $D\alpha$ is the Jacobian of $\alpha$, we can express integration over $W$ as
$\int_{W}f\,{\rm d}V=\int_{U}f\circ\alpha\,\mathrm{V}(D\alpha).$
In the coming proofs, when integrating around a point $x\in\operatorname{Int}\Omega_{i}$, we will change coordinates to the standard basis in $T_{\Omega_{i},x}=\mathbb{R}^{k}$. With this we mean that we can find open sets $W\subset\Omega_{i}$ around $x$ such that the projection map $\pi:W\to B\subset x+T_{\Omega_{i},x}$ is a diffeomorphism, where $x+T_{\Omega_{i},x}\coloneqq\{\,x+y\mid y\in T_{\Omega_{i},x}\,\}$. To integrate over $T_{\Omega_{i},x}$ we use the map $\pi^{-1}$ precomposed with an inclusion map.
More specifically and without loss of generality, by translation and an orthonormal coordinate change, we can assume that $T_{\Omega_{i},x}=\mathbb{R}^{d}\times\{0\}^{n-d}$. In this coordinate system we can write
[…]
where $i$ is the natural inclusion map and $U$ an open subset in $\mathbb{R}^{k}$.
A.2. Important bounds
The following bounds will be used later in our proofs: First, let $T_{\Omega,x},U,W,\pi$ be as in Section A.1. Then for any $y\in W$,
$\left\lVert y-\pi(y)\right\rVert\leq O(\left\lVert x-\pi(y)\right\rVert^{2}).$ (A.2)
This follows since $\Omega_{i}$ is smooth and the tangent space represents the best flat approximation of $\Omega_{i}$ around $x$.
To formulate the second bound, we need the lemma below.
Lemma A.1.
Let $U,W,x,y,\Omega_{i},\pi,i,\alpha$ be as in Section A.1. Then the following holds for the volume function $V$:
$V(D\alpha(y))=1+O(\left\lVert x-\pi(y)\right\rVert^{2}).$
Proof.
Since $\alpha=\pi^{-1}\circ i$, and the tangent space is the best flat approximation of $\Omega$, we can parametrize the $W$ by $\alpha(u)=(u,g(u))$. It is then easy to see that for $i=1,\ldots,d$ we have
$\partial_{i}\alpha(y)=(e_{i},\partial_{i}g(u)),$
where $\partial_{i}g(0)=0$ and $\left\lVert\partial_{i}g(u)\right\rVert=O(\left\lVert u\right\rVert)$. Now
$\det D\alpha_{I}=\begin{cases}1&\text{if $I=(1,2\dots,d)$}\\ O(\left\lVert u\right\rVert)&\text{otherwise.}\end{cases}.$
If we Taylor expand $x\to\sqrt{x}$, we get
$V(D\alpha)=\left(\sum_{I}(\det D\alpha_{I})^{2}\right)^{1/2}=1+O(\left\lVert u\right\rVert^{2}),$
and by applying the above on $(u,0)=x-\pi(y)$ we are finished. ∎
Further, since we have a finite union $\Omega=\cup_{i}\Omega_{i}$, $\Omega_{i}$ is compact, A.2, and the previous lemma implies that we can find a uniform bound $L$ such that for all tuples $(U,W,x,y,\pi,\Omega_{i})$
$\left\lVert y-\pi(y)\right\rVert\leq L\left\lVert x-\pi(y)\right\rVert^{2}$ (A.3)
and
$|V(D\alpha)-1|\leq L\left\lVert x-\pi(y)\right\rVert^{2}$ (A.4)
holds.
A.3. Gamma functions
In the proofs of several of our results we will need to handle the Gamma function $\Gamma(\cdot)$, and both the lower and upper incomplete gamma functions, $\gamma(\cdot,\cdot)$ and $\Gamma(\cdot,\cdot)$ respectively. These are well-known and are defined by the equations
$\Gamma(a)=\int_{0}^{\infty}t^{a-1}e^{-t}\,{\rm d}t,\quad\gamma(a,x)=\int_{0}^{x}t^{a-1}e^{-t}\,{\rm d}t,\quad\Gamma(a,x)=\int_{x}^{\infty}t^{a-1}e^{-t}\,{\rm d}t.$
In this paper both $a$ and $x$ are non-negative real numbers.
We will need the following bounds: First, if $a\geq 1$, then $t^{a-1}\geq x^{a-1}$ and
$\Gamma(a,x)\geq x^{a-1}\int_{x}^{\infty}e^{-t}\,{\rm d}t=x^{a-1}e^{-x}.$ (A.5)
Secondly, if $e^{x}>2^{a}$, then by [20, Theorem 4.4.3],
$\Gamma(a,x)\leq ax^{a-1}e^{-x}.$ (A.6)
Finally, we need the lower bound
$\gamma(a,a)\geq\frac{1}{2}\Gamma(a).$ (A.7)
That this holds can be seen by viewing $\gamma(a,x)$ as an unnormalized version of the cumulative distribution function of the Gamma distribution, for which it is well-known that the median $\nu$ is less than $a$[21].
Appendix B Proof of Theorems 2 and 3
B.1. Proof of Theorem 2
Since $x\to L_{t}^{i}f(x)$ is translation and rotation invariant, we can without loss of generality assume that $\Omega_{i}$ oriented in $\mathbb{R}^{N}$ in such a way which makes it a subset of $\mathbb{R}^{d}\times\{0\}^{N-d}$.
We want to evaluate
$L^{i}_{t}f(x)=\int_{\Omega_{i}}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y.$
We begin by splitting the integral above into
$\int_{\Omega_{i}}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y$ $=\int_{B_{R}(x)\cap\Omega_{i}}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y$
$\quad+\int_{\Omega_{i}\setminus B_{R}(x)}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y$
$=I_{1}+I_{2}.$ (B.1)
For estimating $I_{2}$, by translation invariance we can without loss of generality assume that $x=0$. Now we make a change of variables and rescale $y$, which allows us to say that
$\begin{split}|I_{2}|&=\left|\int_{\Omega_{i}\setminus B_{R}(0)}K_{t}(0,y)(f(0)-f(y))p\,{\rm d}y\right|\\ &=\left|\int_{\Omega_{i}\setminus B_{r_{0}\sqrt{t}}(0)}e^{-\|y\|^{2}/t}v\cdot(-y)p\,{\rm d}y\right|\\ &=\left|\int_{\left(\frac{1}{\sqrt{t}}\Omega_{i}\right)\setminus B_{r_{0}}(0)}e^{-\|y\|^{2}}v\cdot(-y\sqrt{t})t^{d/2}p\,{\rm d}y\right|\\ &\leq t^{d/2+1/2}\int_{\mathbb{R}^{d}\setminus B_{r_{0}}}e^{-\|y\|^{2}}\|y\|p\,{\rm d}y.\end{split}$
Now, by first changing to spherical coordinates and integrating out the angular parts, we deduce that
[…]
To finalize the bound of $I_{2}$, we note that it follows from the assumption $t\leq\frac{R^{2}}{d/2+1}$ that $r_{0}^{2}>\frac{d+1}{2}$, and we can use A.6 and B.2 to conclude
$|I_{2}|\leq B(x)t^{d/2+1/2}e^{-r_{0}^{2}},$ (B.3)
where $B(x)$ is some function such that
$B(x)\leq\frac{d+1}{4}r_{0}^{d}p|\mathbb{S}^{d-1}|.$
To bound $I_{1}$, we use the following simple geometric fact:
$\|x-y\|^{2}=\|\hat{x}-y\|^{2}+\|\hat{x}-x\|^{2}=\|\hat{x}-y\|^{2}+\sin^{2}\theta r^{2}t,$
which implies that
$e^{-\|x-y\|^{2}/t}=e^{-\sin^{2}\theta r^{2}}e^{-\|\hat{x}-y\|^{2}/t}.$
From the above we can conclude
$I_{1}$ $=e^{-\sin^{2}\theta r^{2}}\int_{B_{R}(x)\cap\Omega_{i}}e^{-\|\hat{x}-y\|^{2}/t}v\cdot(x-y)p\,{\rm d}y$
[…]
$=e^{-r^{2}\sin^{2}\theta}(II+III).$ (B.7)
It is easier to integrate over a ball centered around $\hat{x}$, and to this end we define $\delta\geq 0$ by
$\delta=\sqrt{R^{2}-tr^{2}\sin^{2}\theta}.$ (B.8)
Then since $\hat{x}$ is the orthogonal projection of $x$, we have that $B_{R}(x)\cap\Omega_{i}=B_{\delta}(\hat{x})\cap\Omega_{i}$.
Let us focus on $II$: We use the B.8 and change to spherical coordinates, which yields
$II$ $=v\cdot(\hat{x}-x)t^{d/2}\int_{B_{\delta/\sqrt{t}}(\hat{x})\cap\Omega_{i}}e^{-\|\hat{x}-y\|^{2}}p\,{\rm d}y.$
[…]
$=\frac{1}{2}v\cdot\frac{\hat{x}-x}{\|\hat{x}-x\|}t^{d/2+1/2}r\sin\theta|\mathbb{S}^{d-1}|p\gamma(d/2,\delta^{2}/t).$ (B.9)
To estimate the RHS of B.9 we will bound the $\gamma$-function from above and below: Using $r_{0}^{2}\geq\frac{d+2}{2}$, $r<r_{0}/2$ and the definition of $\delta$, we get
$\frac{d}{2}\leq r_{0}^{2}-\sin^{2}\theta r^{2}=\frac{\delta^{2}}{t}.$
By A.7 we now see that
$\frac{1}{2}\Gamma(d/2)\leq\gamma(d/2,d/2)\leq\gamma(d/2,\delta^{2}/t).$ (B.10)
Further, an application of A.5 yields
$\gamma(d/2,\delta^{2}/t)$ $\leq\gamma(d/2,r_{0}^{2})=\Gamma(d/2)-\Gamma(d/2,r_{0}^{2})\leq\Gamma(d/2)-(r_{0}^{2})^{d/2-1}e^{-r_{0}^{2}}$
$=\Gamma(d/2)-r_{0}^{d-2}e^{-r_{0}^{2}}.$ (B.11)
Now B.9, B.10 and B.11 together with $|\mathbb{S}^{d-1}|=\frac{2\pi^{d/2}}{\Gamma(d/2)}$ finally gives
$II=A(d,r_{0},\theta)v_{n,\Omega_{i}}t^{d/2+1/2}r\sin\theta,$
where
$\frac{1}{2}\max(p\pi^{d/2},2\pi^{d/2}p-p\left|\mathbb{S}^{d-1}\right|r_{0}^{d-2}e^{-r_{0}^{2}})\leq A(d,r_{0},\theta)\leq p\pi^{d/2}.$ (B.12)
Finally, $III=0$. This follows from that $B_{R}(x)\cap\partial\Omega_{i}=\emptyset$, the rotational symmetry of $K$, and the fact that the linear function is odd. Collecting B.1, B.3, B.7 and B.9 we get
$L_{t}f(x)=t^{d/2+1/2}\left(A(d,r_{0},\theta)v_{n,\Omega_{i}}\sin\theta re^{-\sin^{2}\theta r^{2}}+B(x)e^{-r_{0}^{2}}\right).$
B.2. Proof of Theorem 3
We will follow the proof of Theorem 2 and modify where needed. Let $I_{2},II$ and $III$ be defined as in B.1 and B.7. Then, since $I_{2}$ is bounded like in B.3, we only need to find bounds for $II$ and $III$.
Let $\delta$ be defined as in B.8 and define $\delta_{0}=\delta/\sqrt{t}$. Recall also the fact that $B_{R}(x)\cap\Omega_{i}=B_{\delta}(\hat{x})\cap\Omega_{i}$. Now the difference in bounding $II$ and $III$ to the proof of Theorem 2 is that $B_{\delta}(\hat{x})\cap\partial\Omega_{i}$ is nonempty. Since, by assumption, $\partial\Omega_{i}$ is part of a $d-1$-dimensional flat space, $B_{\delta}(\hat{x})\cap\Omega_{i}$ is a $d$-dimensional ball, but missing a spherical cap.
We now use cylindrical coordinates $(h,\varrho,\varphi)$ to describe the domain
$B_{\delta/\sqrt{t}}(\hat{x})\cap\Omega_{i}$. In these new coordinates we are centered around $\hat{x}$, and $(\varrho,\varphi)$ are coordinates for a $d-1$-dimensional ball tangential to $\partial\Omega$, while the perpendicular coordinate $h$ is oriented along the outwards normal of $\partial\Omega_{i}$. Let us denote this unit normal by $n_{\partial\Omega}$, and the projection of $\hat{x}$ to $\partial\Omega$ by $\hat{x}_{\partial\Omega}$. We now set $K=(\hat{x}-\hat{x}_{\partial\Omega})\cdot n_{\partial\Omega}=\sqrt{t}k_{0}$, where $-\delta_{0}\leq k_{0}\leq\delta_{0}$.
Then, with $III$ defined in B.7 we get
$III=\int_{-\delta}^{K}\int_{0}^{\sqrt{\delta^{2}-h^{2}}}\int_{\mathbb{S}^{d-2}}K_{t}(\hat{x},y)v\cdot(\hat{x}-y)\varrho^{d-2}\,{\rm d}\varphi\,{\rm d}\varrho\,{\rm d}h.$
We split $v$ into a normal component $v_{n}=(v\cdot n_{\partial\Omega})n_{\partial\Omega}$ and a component $v_{T}=v-v_{n}$ which is tangential to the boundary $\partial\Omega$. Then, since the function $y\to v_{T}\cdot(\hat{x}-y)$ is odd as a function centered around $\hat{x}$, and the domain of integration is symmetric around $\hat{x}$, we know that the tangential component of $III$ satisfies
$III_{T}:=\int_{-\delta}^{K}\int_{0}^{\sqrt{\delta^{2}-h^{2}}}\int_{\mathbb{S}^{d-2}}K_{t}(\hat{x},y)v_{T}\cdot(\hat{x}-y)\varrho^{d-2}\,{\rm d}\varphi\,{\rm d}\varrho\,{\rm d}h=0.$
By definition of $v_{n,\partial\Omega}$, we have that $v_{n}\cdot(\hat{x}-y)=v_{n,\partial\Omega}(n_{\partial\Omega}\cdot(\hat{x}-y))=v_{n,\partial\Omega}h$, which implies that
$III$ $=v_{n,\partial\Omega}\int_{-\delta}^{K}\int_{0}^{\sqrt{\delta^{2}-h^{2}}}\int_{\mathbb{S}^{d-2}}K_{t}(\hat{x},y)h\varrho^{d-2}\,{\rm d}\varphi\,{\rm d}\varrho\,{\rm d}h$
$=v_{n,\partial\Omega}\int_{-\delta}^{K}\int_{0}^{\sqrt{\delta^{2}-h^{2}}}\int_{\mathbb{S}^{d-2}}e^{-h^{2}/t-\varrho^{2}/t}h\varrho^{d-2}\,{\rm d}\varphi\,{\rm d}\varrho\,{\rm d}h$
$=t^{d/2}v_{n,\partial\Omega}\int_{-\delta_{0}}^{k_{0}}he^{-h^{2}}\int_{0}^{\sqrt{\delta_{0}^{2}-h^{2}}}\int_{\mathbb{S}^{d-2}}e^{-\varrho^{2}}\varrho^{d-2}\,{\rm d}\varphi\,{\rm d}\varrho\,{\rm d}h.$
Continuing with the two inner integrals,
$\int_{0}^{\sqrt{\delta_{0}^{2}-h^{2}}}\int_{\mathbb{S}^{d-2}}e^{-\varrho^{2}}\varrho^{d-2}\,{\rm d}\varphi\,{\rm d}\varrho$ $=\frac{|\mathbb{S}^{d-2}|}{2}\int_{0}^{\delta_{0}^{2}-h^{2}}e^{-s}s^{d/2-3/2}\,{\rm d}s$
$=\frac{|\mathbb{S}^{d-2}|}{2}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-h^{2}\right).$
Using this expression in the full integral and applying partial integration in the second equality below yields
$III$ $=t^{d/2}v_{n,\partial\Omega}\frac{|\mathbb{S}^{d-2}|}{2}\int_{-\delta_{0}}^{k_{0}}e^{-h^{2}}h\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-h^{2}\right)\,{\rm d}h$
$=t^{d/2}v_{n,\partial\Omega}\frac{|\mathbb{S}^{d-2}|}{2}\bigg{(}\frac{1}{2}\left[-e^{-h^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-h^{2}\right)\right]^{k_{0}}_{-\delta_{0}}$
$\qquad-\frac{1}{2}e^{-\delta_{0}^{2}}\int_{-\delta_{0}}^{k_{0}}(\delta^{2}-h^{2})^{(d-3)/2}h\,{\rm d}h\bigg{)}$
$=t^{d/2}v_{n,\partial\Omega}\frac{|\mathbb{S}^{d-2}|}{2}\left(\frac{1}{2}e^{-k_{0}^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-k_{0}^{2}\right)+e^{-\delta_{0}^{2}}\frac{(\delta_{0}^{2}-k_{0}^{2})^{(d-1)/2}}{d-1}\right).$
Thus, we know that
$III=t^{d/2}v_{n,\partial\Omega}\frac{|\mathbb{S}^{d-2}|}{2}\left(e^{-\delta_{0}^{2}}\frac{(\delta_{0}^{2}-k_{0}^{2})^{(d-1)/2}}{d-1}+\frac{1}{2}e^{-k_{0}^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-k_{0}^{2}\right)\right).$ (B.13)
We now address the integral $II$ defined in B.7, which means we need to calculate
$J\coloneqq\int_{B_{R}(x)\cap\Omega_{i}}e^{-\|\hat{x}-y\|^{2}/t}p\,{\rm d}y.$
After a change cylindrical coordinates as for $III$, we rewrite this integral as
$J=\int_{-\delta_{0}}^{k_{0}}e^{-h^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-h^{2}\right)\,{\rm d}h.$
We can immediately bound $J$ from above by
[…]
Now we bound $J$ from below: Since the integrand is positive, we can without loss of generality assume that $k_{0}<0$. Then a change of variables $h=-\sqrt{\delta_{0}^{2}-y}$ yields that
$J$ $\geq e^{-\delta_{0}^{2}}\int_{0}^{\delta_{0}^{2}-k_{0}^{2}}e^{y}\gamma\left(\frac{d-1}{2},y\right)\frac{1}{2\sqrt{\delta_{0}^{2}-y}}\,{\rm d}y$
$\geq e^{-\delta_{0}^{2}}\frac{1}{2\delta_{0}}\int_{0}^{\delta_{0}^{2}-k_{0}^{2}}e^{y}\gamma\left(\frac{d-1}{2},y\right)\,{\rm d}y.$
Using partial integration above we then get
$J$ $\geq\frac{e^{-\delta_{0}^{2}}}{2\delta}\left[e^{y}\gamma\left(\frac{d-1}{2},y\right)-\frac{y^{\frac{d-1}{2}}}{\frac{d-1}{2}}\right]_{0}^{\delta_{0}^{2}-k_{0}^{2}}$
$=\frac{e^{-\delta_{0}^{2}}}{2\delta_{0}}\left(e^{\delta_{0}^{2}-k_{0}^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-k_{0}^{2}\right)-\frac{2(\delta_{0}^{2}-k_{0}^{2})^{\frac{d-1}{2}}}{d-1}\right).$
Simplifying further gives us
$J\geq\frac{1}{2\delta_{0}}\left(e^{-k_{0}^{2}}\gamma\left(\frac{d-1}{2},\delta_{0}^{2}-k_{0}^{2}\right)-e^{-\delta_{0}^{2}}\frac{2(\delta_{0}^{2}-k_{0}^{2})^{\frac{d-1}{2}}}{d-1}\right).$ (B.15)
Thus, equation B.13 and the bounds in B.15 and B.14 proves the theorem.
Appendix C Proofs of Section 5
C.1. Proof of Theorem 6
Set $R=2$, let $r_{0}=2/\sqrt{t}$, and under $H_{0}$ $\partial\Omega_{i}\cap B_{4R}(x_{0})=\emptyset$, we can apply Theorem 2 to get an expression for $L_{t}^{i}f(x)$, i.e. for any $x\in B_{1}(x_{0})$ we have that
$L_{t}^{i}f(x)=t^{\frac{d+1}{2}}B(x)e^{-1/t}.$
Furthermore, under our assumption of $H_{0}$, we have that
$|L_{t}^{j}f(x)|=\left\lvert\int_{\Omega_{j}}e^{-\|x-y\|^{2}/t}v\cdot(x-y)p(y)\,{\rm d}y\right\rvert\leq e^{-2/t}$
whenever $j\neq i$. Now applying Theorem 1 we can see that for
$\epsilon=\delta-\max\left\{\sqrt{t}\frac{d+1}{2}\frac{|\mathbb{S}^{d-1}|}{2^{d}|\mathbb{S}^{d}|},1\right\}e^{-2/t},$
thus choosing $t\leq 1$ such that $\sqrt{t}\frac{d+1}{2}\frac{|\mathbb{S}^{d-1}|}{2^{d}|\mathbb{S}^{d}|}e^{-1/t}\leq\delta/2$ we get that
$\mathbb{P}(\exists m:X_{m}\in B_{R}(x_{0});\ |L_{n,t}f(X_{m})|>\epsilon+L_{t}f(X_{m}))\leq 2n\exp\left(-\frac{4e(n-1)\epsilon^{2}}{t}\right)\leq\alpha.$
C.2. Proof of Theorem 7
Under hypothesis $H_{1}^{\prime}$ we have from Theorem 2 that for $r_{0}=1/\sqrt{t}$ that
$L_{t}^{i}f(x)=t^{\frac{d+1}{2}}\left(A(\theta_{i}(x),r_{0},d)v_{n,\Omega_{i}}r_{i}(x)\sin\theta_{i}(x)e^{-r_{i}(x)^{2}\sin^{2}\theta_{i}(x)}+B(x)e^{-r_{0}^{2}}\right),$
whenever $x\in B_{1}(x_{0})$. Also note that $\theta_{i}(x)=0$ if $x\in\Omega_{i}$, see Fig. 3.
The power of rejecting $H_{0}$ when $H_{1}^{\prime}$ is true can be calculated by
$\mathbb{P}(T>\delta).$
Define the events
$A$ :=$:=\{\max_{m}|L_{n,t}f(X_{m})-L_{t}f(X_{m})|<\delta\},$
$B$ :=$:=\{\max_{m}|L_{t}f(X_{m})|\chi_{B_{1}(x_{0})}(X_{m})>2\delta\}.$
Then, for any index $m$ satisfying $B$ we have
$\max_{m}|L_{n,t}f(X_{m})|\geq\max_{m}(|L_{t}f(X_{m})|-|L_{n,t}f(X_{m})-L_{t}f(X_{m})|)\geq\delta,$
which gives together with the union bound that
$\mathbb{P}(T>\delta)$ $\geq\mathbb{P}(A\cap B)$
$=1-\mathbb{P}(A^{c}\cup B^{c})$
$\geq 1-\mathbb{P}(A^{c})-\mathbb{P}(B^{c}).$
We can bound $\mathbb{P}(A^{c})$ immediately using Theorem 1:
$\mathbb{P}(\max_{m}|L_{n,t}f(X_{m})-L_{t}f(X_{m})|>\delta)\leq\alpha.$
We next bound $\mathbb{P}(B^{c})$, using Theorem 2 and the definition of $\delta$:
$\mathbb{P}(B^{c})$ $\leq\mathbb{P}(\max_{m}|L_{t}^{1}f(X_{m})|\chi_{B_{1}(x_{0})\cap\Omega_{2}}(X_{m})<2\delta+L_{t}^{2}f(X_{m})\chi_{B_{1}(x_{0})\cap\Omega_{2}}(X_{m}))$
$\leq\mathbb{P}(\max_{m}|L_{t}^{1}f(X_{m})|\chi_{B_{1}(x_{0})\cap\Omega_{2}}(X_{m})<3\delta)$
$=\mathbb{P}(|L_{t}^{1}f(X_{m})|\chi_{B_{1}(x_{0})\cap\Omega_{2}}(X_{m})<3\delta)^{n}.$
We can now apply Theorem 2 to get that for $x\in B_{1}(x_{0})\cap\Omega_{2}$
$|L_{t}^{1}f(x)|\geq t^{\frac{d+1}{2}}2\pi^{d/2}|v_{n,\Omega_{1}}|\frac{\|x-x_{0}\|}{\sqrt{t}}|\sin\theta_{1}(x)|e^{-\|x-x_{0}\|^{2}\sin^{2}\theta_{1}(x)/t}-\delta.$
Denote $G(x)=xe^{-x^{2}}$, then $G^{-1}(y)=\sqrt{-W_{0}(-2y^{2})/2}$ and $G^{-1}(y)=\sqrt{-W_{-1}(-2y^{2})/2}$ are the two solutions to $G(x)=y$, and $W$ is the Lambert $W$ function and the subindex denotes the two different branches. Thus, we can estimate
$\mathbb{P}(|L_{t}^{1}f(X_{m})|\chi_{B_{1}(x_{0})\cap\Omega_{2}}(X_{m})<3\delta)\leq\mathbb{P}(G(\|X_{m}-x_{0}\|\sin\theta_{1}/\sqrt{t})\chi_{B_{1}(x_{0})\cap\Omega_{2}}<\frac{\delta}{t^{\frac{d+1}{2}}C})$
where $C=\pi^{d/2}|v_{n,\Omega_{1}}|/2$. Using the inverse of $G$ as above we get the following terms
$P_{1}$ $:=\mathbb{P}\left(\|X_{m}-x_{0}\|<\frac{\sqrt{-tW_{0}\left(-2\frac{\delta^{2}}{t^{d+1}C^{2}}\right)}}{\sqrt{2}\sin\theta_{1}};X_{m}\in B_{1}(x_{0})\cap\Omega_{2}\right),$
$P_{2}$ $:=\mathbb{P}\left(\|X_{m}-x_{0}\|>\frac{\sqrt{-tW_{-1}\left(-2\frac{\delta^{2}}{t^{d+1}C^{2}}\right)}}{\sqrt{2}\sin\theta_{1}};X_{m}\in B_{1}(x_{0})\cap\Omega_{2}\right),$
$P_{3}$ $:=\mathbb{P}(X_{m}\in(B_{1}(x_{0})\cap\Omega_{2})^{c}).$
For $P_{1}$ we note that $W_{0}$ satisfies
$-W_{0}(-\varrho)<e\varrho,$
for $0\leq\varrho\leq 1$, as such
$P_{1}\leq\mathbb{P}\left(\|X_{m}-x_{0}\|<\frac{e\delta}{t^{d/2}C\sin\theta_{1}};X_{m}\in\Omega_{2}\right)\leq|\Omega_{2}|\left(\frac{e\delta}{t^{d/2}C\sin\theta_{1}}\right)^{d}.$
Thus, if we require that
$\frac{e\delta}{t^{d/2}C\sin\theta_{1}}\leq 1/4,$
we have that $P_{1}<\mathbb{P}(X_{m}\in B_{1/4}(x_{0}))$. Secondly, we know that
$-W_{-1}(-\varrho)>\log 1/\varrho,$
for $10<\varrho<$ 1, and as such
$P_{2}\leq\mathbb{P}\left(\|X_{m}-x_{0}\|>\frac{\sqrt{t\log\left(\frac{t^{d+1}C^{2}}{2\delta^{2}}\right)}}{\sqrt{2}\sin\theta_{1}};X_{m}\in B_{1}(x_{0})\cap\Omega_{2}\right).$
Which if
$t\left((d+1)\log(1/t)+\log\left(\frac{2^{3}}{C^{2}}\right)\right)\leq 2(1-\sin^{2}\theta_{1}),$
is bounded as
$P_{2}\leq\mathbb{P}(X_{m}\in(B_{1}(x_{0})\setminus B_{1/2}(x_{0}))\cap\Omega_{2}).$
In conclusion, we have that for large enough $n$, $P_{1}+P_{2}+P_{3}<1$. Assembling the above we have proven our theorem.
Appendix D Proof of Theorem 4 and Theorem 5
D.1. Proof of Theorem 4
We begin by splitting up the domain $\Omega_{i}$:
$\begin{split}L^{i}_{t}f(x)&=\int_{\Omega_{i}}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y\\ &=\int_{\Omega_{i}\cap B_{1}(x)}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y\\ &\quad+\int_{\Omega_{i}\setminus B_{1}(x)}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y\\ &=I+II.\end{split}$ (D.1)
We first note that
$II=\int_{\Omega_{i}\setminus B_{R}(x)}K_{t}(x,y)(f(x)-f(y))p\,{\rm d}y\leq e^{-R^{2}/t}\operatorname{diam}(\Omega).$ (D.2)
To estimate $I$ we will make a change of variables to the tangent space at $x_{0}$ and use arguments similar to those in the proof of Theorem 2. Specifically, let $\pi:\Omega_{i}\cap B_{R}(x)\to T_{\Omega_{i},_{x}}\cap B_{R}(x)$ be the projection map, and $\alpha=\pi^{-1}\circ i:\mathbb{R}^{d}\cap B_{R}(0)\to\Omega_{i}\cap B_{R}(x)$ a coordinate chart as in (A.1). We will use $\alpha$ to integrate over $T_{\Omega_{i},x_{0}}$.
To simplify notation, we will use $\hat{x}$ and $\hat{y}$ to denote both $\pi(x),\pi(y)\in\mathbb{R}^{N}$, and sometimes implicitly assume the projection $i^{-1}$ such that $\hat{x},\hat{y}\in\mathbb{R}^{d}$. The space in which these points lie should be clear from context.
Before making the coordinate change, we find bounds relating $K(x,y)$ to $K(x,\hat{y})$: We recall that $K_{t}(x,y)=e^{\frac{-\left\lVert x-y\right\rVert^{2}}{t}}$, and by elementary considerations we get
$e^{-\frac{\left\lVert x-\hat{y}\right\rVert^{2}}{t}-3\frac{\left\lVert y-\hat{y}\right\rVert^{2}}{t}}\leq e^{-\frac{\left\lVert x-y\right\rVert^{2}}{t}}\leq e^{-\frac{\left\lVert x-\hat{y}\right\rVert^{2}}{t}+3\frac{\left\lVert y-\hat{y}\right\rVert^{2}}{t}}.$ (D.3)
Since $\Omega_{i}$ is $(L,2R)$-regular, we use (A.3) and the fact that $y\in B_{2R}(x)$ to conclude
$\left\lVert y-\hat{y}\right\rVert\leq L\left\lVert x_{0}-\hat{y}\right\rVert^{2}\leq L4R^{2},$
which together with (D.3) yields
$e^{-3(L4R^{2})^{2}/t}K_{t}(x,\hat{y})\leq K_{t}(x,y)\leq e^{3(L4R^{2})^{2}/t}K_{t}(x,\hat{y}).$
Furthermore, since $(L4R^{2})^{2}/t\leq\frac{1}{4}$ by our assumptions, we have the bounds
$e^{3(L4R^{2})^{2}/t}\leq 1+4(L4R^{2})^{2}/t\quad\mathrm{and}\quad e^{-3(L4R^{2})^{2}}\geq 1-3(L4R^{2})^{2}/t.$
Thus,
$|K_{t}(x,y)-K_{t}(x,\hat{y})|\leq 4\frac{(L4R^{2})^{2}}{t}K_{t}(x,\hat{y}).$ (D.4)
Replacing $K_{t}(x,y)$ with $K_{t}(x,\hat{y})$ in $I$ we get
$I=\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(x,\hat{y})(f(x)-f(y))p\,{\rm d}y+E_{1},$ (D.5)
and using D.4 it holds that
$|E_{1}|\leq 4\frac{(L4R^{2})^{2}}{t}R\left|\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(x,\hat{y})p\,{\rm d}y\right|.$ (D.6)
We now decompose the integral of the first term in D.5 as follows
$\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(x,\hat{y})(f(x)-f(y))p\,{\rm d}y=\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(x,\hat{y})(f(x)-f(\hat{y}))p\,{\rm d}y$
$\qquad+\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(x,\hat{y})(f(\hat{y})-f(y))p\,{\rm d}y=I_{1}+I_{2}$ (D.7)
The quantity $I_{2}$ will be treated like an error term. Using A.3 we see that
$|I_{2}|\leq\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(x,\hat{y})L\left\lVert\hat{y}-x_{0}\right\rVert^{2}p\,{\rm d}y.$
Now we make a coordinate change with $\alpha$ and use the bound on the volume form in A.4 to get
$\begin{split}&\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(x,\hat{y})L\left\lVert\hat{y}-x_{0}\right\rVert^{2}p\,{\rm d}y\\ &\qquad\leq LR^{2}\int_{T_{\Omega_{i},x_{0}}\cap B_{R}(x)}K_{t}(x,\hat{y})(1+L\left\lVert x_{0}-\hat{y}\right\rVert^{2})p\,{\rm d}\hat{y}\\ &\qquad\leq LR^{2}(1+L4R^{2})\int_{T_{\Omega_{i},x_{0}}\cap B_{R}(x)}K_{t}(x,\hat{y})p\,{\rm d}\hat{y}.\end{split}$
The RHS of the above display can be handled similarly to B.9, which means
$\begin{split}|I_{2}|&\leq LR^{2}(1+L4R^{2})\left|\mathbb{S}^{d-1}\right|t^{d/2}p\Gamma(d/2)=LR^{2}(1+L4R^{2})t^{d/2}2p\pi^{d/2}.\end{split}$
We proceed now with $I_{1}$ from Section D.1, which we want to estimate as accurately as possible. Using the coordinate change $\alpha$ and A.4 we write
$I_{1}$ $=e^{-r^{2}\sin^{2}\theta}\int_{\Omega_{i}\cap B_{R}(x)}K_{t}(\hat{x},\hat{y})(f(x)-f(\hat{y}))p\,{\rm d}y$
$=e^{-r^{2}\sin^{2}\theta}C_{1}(x)\int_{T_{\Omega_{i},x_{0}}\cap B_{R}(x)}K_{t}(\hat{x},\hat{y})(f(x)-f(\hat{y}))p\,{\rm d}\hat{y},$ (D.8)
where $C_{1}(x)>0$ is such that $|C_{1}-1|\leq(1+L4R^{2})$.
The integral on the right in Section D.1 is exactly $II$ from B.7, which we compute as in B.9:
$\int_{T_{\Omega_{i},x_{0}}\cap B_{R}(x)}K_{t}(\hat{x},\hat{y})(f(x)-f(\hat{y}))p\,{\rm d}\hat{y}=A(d,r_{0}\theta)v_{\Omega_{i}}t^{d/2+1/2}r\sin\theta,$ (D.9)
where $A(d,r_{0},\theta)$ is as in (B.12). Now, from Sections D.1, D.1 and D.9 we have
$\begin{split}I_{1}+I_{2}=C_{1}(x)A(d,r_{0},\theta)v_{\Omega_{i}}t^{d/2+1/2}r\sin\theta e^{-r^{2}\sin^{2}(\theta)}+C_{2}(x)LR^{2}(1+L4R^{2})t^{d/2}2p\pi^{d/2},\end{split}$
for a function $|C_{2}(x)|\leq 1$. This combined with the split in (D.5) and (D.6) gives us for another function $|C_{3}(x)|\leq 1$
$I=$ $I_{1}+I_{2}+E_{1}=C_{1}(x)A(d,r_{0},\theta)v_{\Omega_{i}}t^{d/2+1/2}r\sin\theta e^{-r^{2}\sin^{2}(\theta)}$
$+C_{3}(x)C_{L,R}t^{d/2}2p\pi^{d/2}$
where
$C_{L,R}:=4\frac{(L4R)^{4}}{t}R+LR^{2}(1+L4R^{2}).$
Finally, the above and (D.2) finishes the proof.
D.2. Proof of Lemma 4.2
First applying Theorem 4 $L_{t}^{i}f(x)$, and then using the $(L,2R)$ regularity of $\Omega_{i}$, we use A.3 to get
$|r\sin\theta|=\frac{\|x-\hat{x}\|}{\sqrt{t}}\leq\frac{LR^{2}}{\sqrt{t}}<\frac{1}{2^{4}}.$
As such we have that $e^{-2^{-8}}|r\sin\theta|\leq|r\sin\theta|e^{-r^{2}\sin^{2}\theta}\leq|r\sin\theta|$. Now applying Theorem 4 the conclusion follows.
D.3. Proof of Theorem 5
To simplify notation, let $h_{j}=x-X_{j}$.
$\mathbb{E}_{\epsilon}L_{n,t}f(x)$ $=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\epsilon}K_{t}(x,X_{j})(f(x)-f(X_{j}+\epsilon_{j}))$
$=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\epsilon}e^{-\left\lVert h_{j}-\epsilon_{j}\right\rVert^{2}/t}v\cdot(h_{j}-\epsilon_{j})$ (D.10)
Let us compute a single term in the sum in Section D.3: Since the expectation is w.r.t $\epsilon$ we can treat $h=h_{j}$ as fixed, and then algebraic manipulations give us for $z\sim\mathcal{N}(0,\sigma^{2}I)$
$\mathbb{E}_{z}e^{-\left\lVert h+z\right\rVert^{2}/t}(h+z)$
$\quad=(2\pi\sigma^{2})^{-N/2}\int_{\mathbb{R}^{N}}e^{-\left\lVert h+z\right\rVert^{2}/t}e^{-\left\lVert z\right\rVert^{2}/(2\sigma^{2})}v\cdot(h+z)\,{\rm d}z$
$\quad=(2\pi\sigma^{2})^{-N/2}\int_{\mathbb{R}^{N}}e^{-\left(\left\lVert h\right\rVert^{2}/t+2\langle h,z\rangle/t+\left\lVert z\right\rVert^{2}/t+\left\lVert z\right\rVert^{2}/(2\sigma^{2})\right)}v\cdot(h+z)\,{\rm d}z$
$\quad=(2\pi\sigma^{2})^{-N/2}e^{-\frac{\left\lVert h\right\rVert^{2}}{2\sigma^{2}+t}}\int_{\mathbb{R}^{N}}e^{-\frac{1}{kt}\left\lVert z+kh\right\rVert^{2}}v\cdot(h+z)\,{\rm d}z.$
In the second to last step we completed the square and used $k=\frac{2\sigma^{2}}{2\sigma^{2}+t}$. This last integral can be viewed as the expectation
$(\pi kt)^{-N/2}\int_{\mathbb{R}^{N}}e^{-\frac{1}{kt}\left\lVert z+kh\right\rVert^{2}}v\cdot(h+z)\,{\rm d}z=\mathbb{E}_{X}[(h+X)]=(1-k)v\cdot h$
where $X\sim N(-kh,I\frac{kt}{2})$. Then we can conclude that
$\mathbb{E}_{z}e^{-\left\lVert h+z\right\rVert^{2}/t}v\cdot(h+z)=v\cdot h\frac{t^{N/2+1}}{(2\sigma^{2}+t)^{N/2+1}}e^{-\frac{\left\lVert h\right\rVert^{2}}{2\sigma^{2}+t}}.$
Appendix E Proof of Theorem 1
First, using the union bound we get
$\begin{split}&P\bigg{(}\max_{i}\left|L_{n,t}f(X_{i})-\frac{n-1}{n}L_{t}f(X_{i})\right|>\epsilon\bigg{)}\\ &\qquad\leq\sum_{i=1}^{n}P\left(\left|L_{n,t}f(X_{i})-\frac{n-1}{n}L_{t}f(X_{i})\right|>\epsilon\right).\end{split}$ (E.1)
Using the definitions of $L_{n,t}$ and $L_{t}$, see 3.1 and 3.2, and using that the random variables $X_{1},\dots,X_{N}$ are i.i.d., we can replace each $X_{i}$ by $X_{1}$ in each term in the summand of E.1. Let $Z$ be an independent copy of $X_{1}$. Then each summand in E.1 equals
$\begin{split}P\bigg{(}&\bigg{|}\frac{1}{n}\sum_{j=1}^{n}K_{t}(X_{1},X_{j})(f(X_{1})-f(X_{j}))\\ &\quad-\frac{n-1}{n}\mathbb{E}_{Z}[K_{t}(X_{1},Z)(f(X_{1})-f(Z))]\bigg{|}>\epsilon\bigg{)}.\end{split}$ (E.2)
To simplify notation, we denote
$W_{i}(x)=K_{t}(x,X_{i})(f(x)-f(X_{i}))\quad\mathrm{and}\quad Y_{i}(x)=W_{i}(x)-\mathbb{E}_{X_{i}}[W_{i}(x)].$
We now rewrite E.2 as
$P\left(\left|\frac{1}{n-1}\sum_{i=2}^{n}Y_{i}(X_{1})\right|>\frac{n}{n-1}\epsilon\right).$
Now by the tower property we have that
$P\left(\left|\frac{1}{n-1}\sum_{i=2}^{n}Y_{i}(X_{1})\right|>\frac{n}{n-1}\epsilon\right)=\mathbb{E}\left[P\left(\left|\frac{1}{n-1}\sum_{i=2}^{n}Y_{i}(X_{1})\right|>\frac{n}{n-1}\epsilon\mid X_{1}\right)\right]$
In order to use Hoeffding’s inequality we need to show that $Y_{i}(x)$ is a bounded random variable for all $x\in\Omega$. First,
$W_{i}(x)$ $=K_{t}(x,X_{i})(f(x)-f(X_{i}))=e^{-\left\lVert x-X_{i}\right\rVert^{2}/t}v\cdot(x-X_{i})\leq e^{-\left\lVert x-X_{i}\right\rVert^{2}/t}\|x-X_{i}\|$
$\leq\sup_{\rho}e^{-\rho^{2}}\rho\sqrt{t}\leq\sqrt{\frac{t}{2e}}.$
Now Hoeffdings inequality states that (where $C_{n}=\frac{n}{n-1}$)
$\mathbb{P}\left(\left|\frac{1}{n-1}\sum_{i=2}^{n}Y_{i}\right|>C_{n}\epsilon\mid X_{1}\right)\leq 2\exp\left(-\frac{4e(n-1)C_{n}^{2}\epsilon^{2}}{t}\right)$
and the proof is complete after taking expectations.
Appendix F Construction of the neural network dataset
We will limit ourselves to spherical neural nets with 3 nodes, i.e. $W,W^{*}\in\mathbb{R}^{6}$. We will also set $\mathbf{B}=[-10,10]^{6}$.
Given a set of data-points $\mathcal{D}=\{x_{1},x_{2},\ldots,x_{n}\}$ we are interested in sets of the form
$\widehat{\Omega}_{\delta}:=\{W\in\mathbf{B}:|f_{W}(x_{i})-f_{W^{\ast}}(x_{i})|<\delta,\quad(x_{i},f_{W^{\ast}}(x_{i}))\in\mathcal{D}\}.$
To try to limit computations, we use interval constraint propagation, and more specifically forward-backward propagation in the computational tree of the expressions $|f_{W}(x_{i})-f_{W}^{*}(x_{i})|<\delta$, where $\delta\approx 10^{-16}$, or machine epsilon, and the set of samples $\mathcal{D}$ has size 100 to produce a paving (a covering of boxes). For more information about these methods, see [22, 23, 24]. The result of this paving is a set of boxes $\mathbb{X}^{+}$, where each box in $\mathbb{X}^{+}$ has a maximum width of $0.01$.
To construct a point of the zero set we will simply choose the centroid of each box as a representative. By construction, we know that such a centroid is at most $0.005\sqrt{6}$ distance away from a point in $\widehat{\Omega}_{\delta}$.